A sinusoidal transverse wave traveling in the positive direction of an x-axis has amplitude of 2.0cm, a wavelength of 10c and a frequency of 400Hz. If the wave equation is of the form
1
Expert's answer
2020-03-23T10:59:21-0400

As per the given question,

Amplitude(A)= 2cm

Wavelength "(\lambda)" =10cm

Frequency (f)=400Hz

We know that general equation of wave

"y=A\sin(\omega t+\phi)"

"\Rightarrow \omega = 2\pi f=2\pi \times 400=800\pi"

Phase angle "\phi=0"

Now, substituting the values

"y=2 \sin(800\pi t+0)"

"\Rightarrow y= 2\sin (800 \pi t)"